I need help with step by step instructions that show how to solve:
x/2+2y/5=6
x/2-3y/5=-14

Answer by RAY100(1637)

(1) x/2 +2y/5 =6
.
mult by lcd = 10,,,,,5x +4y = 60,,,,,(1a)
.
(2) x/2 -3y/5 = -14
.
mult by lcd = 10,,,,,,5x -6y = -140,,,(2a)
.
(1a) -(2a) = 10y = 200,,,,,y= 20,,,,,,,,,
.
subst in (1a) 5x +4(20) =60
.
5x +80 =60
.
5x = -20
,
x=-4,,,,
.
check,,,,,,(-4,20),,,(1) (-4)/2 +2(20)/5 =-2 + 8 =6,,,,ok
.
(2) (-4)/2 -3(20)/5 = -2 -12 = -14,,,,,ok
.